Suppose that x and y vary inversely and x=4 when y=7 Write the function that models the inverse variation
a. y+1.75 b. y= 28/x c. y=11/x d. y=3/x

Answer by stanbon(48569)

Suppose that x and y vary inversely
x = k/y
-------------------
and x=4 when y=7
So, solve for "k":
4 = k/7
k=28
-----------
Now you know the equation is:
x = 28/y
or y = 28/x
This is the function that models the inverse variation
-------------------------
a. y+1.75 b. y= 28/x c. y=11/x d. y=3/x
-------------
Answer: b
===========================
Cheers,
Stan H.